# Generative Psycho-Lexical Approach for Constructing Value Systems in Large Language Models

### v1

@@ 5.2 LLM Safety Prediction for Evaluating Predictive Validity
Predictive validity refers to how well measurement results based on a particular value system can anticipate future decisions, actions, or outcomes [7]. In the context of human values, evaluating predictive validity often involves examining the correlations between values and other psychological constructs, such as personality traits and attitudes. For LLMs, we propose to evaluate the predictive validity of a value system by examining the relationship between its values and LLM safety, arguably one of the most critical concerns and "psychological constructs" for LLM deployment.
To this end, we draw prompts and safety scores from the established safety benchmark SALAD-Bench [47]. We administer the prompts to LLMs, collect their responses, and measure the revealed values using GPV [92]. We follow the linear probing protocol [15] to evaluate the predictive validity of the value representations under different value systems. Using the Bradley-Terry model [11], we train a linear classifier with pairwise cross-entropy loss to predict the relative safety of LLM pairs:
@@
Here, $s_{M_{i}}\in\mathbb{R}$ is the predicted LLM safety score, and $\mathbf{x}_{M_{i}}\in\mathbb{R}^{|V_{H}|}$ is the vector of LLM values under a particular value system. In this case, we only care about the relative safety of LLM pairs, not absolute safety scores. Therefore, the predictive validity is indicated by the binary classification accuracy of a well-trained linear classifier.
@@ 5.3 LLM Value Alignment for Evaluating Representation Power
Values are cognitive representations of motivational goals [76]. A well-constructed value system should be able to fully represent the broad motivations behind LLM outputs, therefore enabling effective transituational alignment of LLMs with human values.
Our evaluation builds on the BaseAlign algorithm [89], originally designed to align LLM outputs with Schwartz’s basic human values. This approach employs a target human value vector, either heuristically customized or derived from human survey data. We extend BaseAlign to arbitrary value systems by distilling the target vector from the human preference:
@@
Here, $\mathbf{x}_{V}^{*}\in\mathbb{R}^{|V|}$ is the target value vector for values $V$ under evaluation. It is distilled from dataset $\mathcal{D}$ [59] using open-vocabulary value evaluator $M_{E}$ [92]. $\mathcal{D}$ contains triplets of prompt $p$, winning response $r_{w}$, and losing response $r_{l}$. The distillation objective is to minimize the L1 distance between the target value and the values of winning responses, while simultaneously maximizing the distance with those of losing responses. Given that the objective function is piecewise linear, the optimal solution can be derived by iteratively examining each value dimension $V_{i}$ and the corresponding value measurements $\mathcal{M}_{V,i}=\{M_{E}(p,r_{w},V_{i})\}_{\mathcal{D}}\cup\{M_{E}(p,r_{l},V_{i})\}_{\mathcal{D}}$:
@@
Next, we adopt PPO [73] to align the LLM with the target value under $V$ [89]:
$\min_{\theta}\mathbb{E}_{p\sim\mathcal{D}_{p},r\sim\pi_{\theta}(p)}\left|\mathbf{x}_{V}^{*}-M_{E}(p,r,V)\right|,$ (5)
where $\theta$ is the LLM policy parameters, $\mathcal{D}_{p}$ is the dataset of prompts [17], and $\pi_{\theta}(p)$ is the parameterized LLM response distribution. The representation power of $V$ is evaluated by the helpfulness and harmlessness of the aligned LLMs [5, 89].
@@ 6 Results
This section presents the results of our experiments, including the proposed LLM value system, the analysis of the system, and the benchmarking of our system against the well-established Schwartz’s values [76].
@@ 6.1 Proposed Value System
Fig. 3 visualizes the value system constructed by GPLA. Table 1 gathers its factor loadings, Cronbach’s Alpha, and confidence intervals. The factor loadings indicate the strength of the relationship between each factor and its atomic values, with higher loadings suggesting more contribution to the factor. Cronbach’s Alpha measures the internal consistency of each factor, with higher values indicating greater reliability. Some atomic values are removed to ensure clear loading patterns and desirable factor reliability [1]. After that, all our factors reach the standard psychometric threshold of 0.7, indicating strong internal consistency [85].
@@
Factor Value Loading Cronbach’s Alpha (CI)
Social Responsibility Equity 0.890 0.957 (0.952 – 0.961)
@@ 6.2 Analyzing Value System
Fig. 3 (Top) presents the correlations between the factors in our value system. Similar to Schwartz’s theory of basic human values [76], LLM values also exhibit both compatible and opposing relationships. Notably, social responsibility, rule-following, and rationality show positive correlations with one another, while all of them are negatively correlated with risk-taking.
@@
We conduct circumplex analysis based on the correlations between factors. Circumplex analysis is a statistical method that examines whether the underlying structure of variables aligns with a circumplex pattern, and, if so, the positions of variables on a circle. The stronger the correlation between variables, the shorter their distance on the circumference. Fig. 3 (Bottom) illustrates the analysis results based on Browne’s circular stochastic process model [12, 28]. The compatible values are closer on the circle (e.g., Social Responsibility and Rule-Following) while opposing values are positioned diagonally (e.g., Risk-Taking and Rule-Following). The results verify the presence of a circumplex structure in the value system.
@@
To evaluate the consistency and robustness of our multivariate system structure (i.e., the 5-dimensional relationship), we measure LLM values using two distinct prompt datasets: the psychometric prompts from GPV [92] and the red-teaming prompts from SALAD-Bench [47], which feature distinct prompt distributions. Their measurements yield an average intra-LLM correlation of 0.87; here we use intra-LLM correlations because the relative value hierarchy within an LLM is more important than their absolute measurements. This result indicates a high level of consistency in the value structure across prompt distributions. We also find a high correlation (0.73) between intra-LLM value consistency and LLM safety scores [47]. It suggests that LLMs with higher value consistency tend to be safer. Complete results are available in § B.3.
@@ 6.3 Comparing Value Systems
We benchmark the proposed value system against Schwartz’s value system [76], the most established framework for human values and commonly used in LLM value studies.
@@
We follow the standard validation procedures of CFA [79] to evaluate the structure validity of different value systems. For our value system, we construct it using half of the measurement data and bootstrap the data to ensure its sufficiency (#data points 5$\geq$ 5$\times$ 5 × #variables). The other half of the data is held out for CFA. For Schwartz’s value system, we map the observed value variables (the atomic values in our system) to its four high-level values or ten low-level values, according to the semantic relevance [79] measured by an embedding model [58]; all data is used for CFA. Table 2 displays the CFA results. Our value system demonstrates a better fit for the data, capturing the underlying values behind LLM generations.
Value system #Values CFI $\uparrow$ GFI $\uparrow$ RMSEA $\downarrow$ AIC $\downarrow$ BIC $\downarrow$
Schwartz (H) 4 0.56 0.52 0.10 340 1484
@@
Table 2: CFA results for different value systems. H and L denote high and low-level Schwartz values, respectively.
@@
We follow the experimental setup in [92, Section 5.2] to predict LLM safety based on their value orientations. Table 3 presents the prediction accuracy under different value systems. The higher accuracy of our value system indicates its superior predictive validity. In addition, according to the parameters of well-trained linear classifiers, we find that Social Responsibility, Rule-Following, and Rationality enhance safety, whereas Risk-taking and Self-Competence undermine it; see § B.5 for details.
Value system Acc (%)
Schwartz (H) 81$\pm$ 15
Schwartz (L) 74$\pm$ 16
Ours 87$\pm$ 9
Table 3: Accuracy of LLM safety prediction based on values.
@@
We follow the experimental setup in [89, Section 6.2] to perform LLM value alignment. Different value systems are respectively used to represent LLM outputs and desired human values. We employ GPV [92] as an open-vocabulary value evaluator for all value systems, but also include the original results of [89] using its Schwartz-specific evaluator for comparison. Table 4 shows the alignment performance of different value systems. Alignment under our value system converges to the lowest harmlessness and the highest helpfulness, establishing its superior representation power. Full experimental details are available in § A.4.
Model Harmlessness $\downarrow$ Helpfulness $\uparrow$
Schwartz* [89] -1.52 2.15
@@
Table 4: Alignment performance of different value systems. Schwartz* denotes the original results drawn from [89] using a Schwartz-specific value evaluator. Both Schwartz baselines are based on the 10-dimensional Schwartz values [89].
@@ 7 Conclusion
This study presents the Generative Psycho-Lexical Approach (GPLA), a novel paradigm for constructing psychologically grounded value systems. GPLA enables automated, scalable, adaptable, and non-reactive value system construction. With GPLA, we introduce the first theoretically and empirically validated five-factor LLM value system, encompassing Social Responsibility, Risk-Taking, Rule-Following, Self-Competence, and Rationality. We accompany the value system with three benchmarking tasks, rooted in both psychological theories and AI priorities. Experimental results confirm the superior validity and utility of the proposed value system compared to the canonical Schwartz’s values.
Our findings underscore the potential of GPLA to address key limitations of traditional psycho-lexical methods, by offering adaptability to evolving contexts and agent-specific requirements. The proposed value system paves the way for more effective evaluation, understanding, and alignment of LLMs, thereby contributing to their safe and reliable deployment. Future work will extend GPLA to multilingual and scenario-specific contexts, and apply the proposed value system for further LLM evaluation and alignment.
@@
This study is limited to identifying and measuring values in English. Values are both culturally and linguistically sensitive [77], and LLM value orientations can vary across languages [13]. Therefore, future research should extend GPLA to multilingual contexts.
In addition, we follow BaseAlign [89] when benchmarking value systems on value alignment. However, the algorithm, like many existing alignment methods, assumes a single, consistent alignment target. Future work should explore distributional and pluralistic alignment [83, 96] to more holistically evaluate a value system.
@@ A.1 Data Statistics for Collecting Value Lexicons
We collect value-laden LLM generations from four data sources: ValueBench [65], GPV [92], ValueLex [9], and BeaverTails [35]. They provide data of different forms: raw LLM responses, parsed perceptions (a sentence that is highly reflective of values [92]), and annotated values. The summary of the data statistics is shown in Table 5.
ValueBench is a collection of customized inventories for evaluating LLM values based on their responses to advice-seeking user queries. By administering the inventories to a set of LLMs, the authors collect 11,928 responses [^1] https://github.com/Value4AI/ValueBench/blob/main/assets/QA-dataset-answers-rating.xlsx, each considered as one perception. The responses are annotated with 37,526 values by Kaleido [82], of which 330 are unique.
GPV [92] is a psychologically grounded framework for measuring LLM values given their free-form outputs. Perceptions are considered atomic measurement units in GPV, and the authors collect 24,179 perceptions [^2] https://github.com/Value4AI/gpv/blob/master/assets/question-answer-perception.csv from a set of LLM subjects. The perceptions are annotated with 62,762 values, of which 361 are unique.
In ValueLex [9], the authors collect 745 unique values from a set of fine-tuned LLMs via direct prompting (see Appendix C for more details).
BeaverTails [35] is an AI safety-focused collection. We use a subset of the BeaverTails dataset [^3] https://huggingface.co/datasets/PKU-Alignment/BeaverTails/tree/main/round0/30k, which contains 3012 LLM responses, which are then parsed into 10,008 perceptions. The perceptions are annotated with 21,968 values, of which 395 are unique.
@@
Source #perceptions #values #unique values
ValueBench 11,928 37,526 330
@@ A.2 LLM Subjects
Our experiments involve 33 LLMs coupled with 21 profiling prompts [67]. The LLMs and profiling prompts are listed in Table 6 and Table 7, respectively.
Model #Params
Baichuan2-13B-Chat 13B
@@
No. Value Anchoring Prompt
1 Default You are a helpful assistant.
@@
Table 7: Anchoring prompts for LLM profiling [67].
@@ A.3 Value Measurement for Structuring Value System
We measure the value orientations of the LLMs subjects following GPV [92]. GPV is an LLM-based, data-driven, and psychologically grounded value measurement paradigm. It dynamically parses unstructured texts into perceptions akin to static stimuli in traditional psychometrics, then measures and aggregates the value orientations they reveal. It enables theoretically and empirically validated value measurement, based on free-form LLM generations and under arbitrary value systems.
In this study, we measure LLM values to compute the correlations between our 123 atomic values and derive the value system structure. Similar to [92], we generate one advice-seeking and value-eliciting prompt for each of the atomic values, using LABEL:prompt:item_generation. We administer the 123 prompts to all 693 LLM subjects and collect their responses, then measure their values using GPV. Each value dimension is measured using all responses. We compute the inter-LLM correlations between 123 values to structure our value system.
@@ A.4 LLM Value Alignment
All experiments were conducted on two NVIDIA L20 GPUs, each with 48GB of memory. We generally follow the experimental setup in [89], with the exceptions noted below. As shown in Table 4, our modifications improve the harmlessness of the aligned model with only marginal reduction in helpfulness.
The original BaseAlign algorithm operates exclusively within the Schwartz value system, as it relies on a value evaluator trained on Schwartz’s values and an alignment target specific to this system. We extend BaseAlign to align LLMs under any arbitrary value system. First, we employ GPV [92] as an open-vocabulary value evaluator. Second, we propose a method for distilling the alignment target from human preference data (§ 5.3). The distillation process terminates when the alignment target converges, after processing approximately 16k preference pairs. The results of this distillation are shown in Table 8 for Schwartz’s values and Table 9 for our system. The distilled alignment target, based on Schwartz’s values, closely matches the heuristically defined target in BaseAlign [89], demonstrating the effectiveness of our approach.
@@
Value Original target Distilled target
Self-Direction 0.0 0.0
@@
Table 8: Alignment targets for Schwartz’s values, on a scale from -1 to 1. Original: heuristically defined target in BaseAlign [89]. Distilled: distilled target from human preference data.
Value Target
Social Responsibility 1.0

### v2

@@ 5.2 LLM Safety Prediction for Evaluating Predictive Validity
Predictive validity refers to how well measurement results based on a particular value system can anticipate future decisions, actions, or outcomes bardi2003values. In the context of human values, evaluating predictive validity often involves examining the correlations between values and other psychological constructs, such as personality traits and attitudes. For LLMs, we propose to evaluate the predictive validity of a value system by examining the relationship between its values and LLM safety, arguably one of the most critical concerns and "psychological constructs" for LLM deployment.
To this end, we draw prompts and safety scores from the established safety benchmark SALAD-Bench li2024salad. We administer the prompts to LLMs, collect their responses, and measure the revealed values using GPV ye2025gpv. We follow the linear probing protocol chen2020simple to evaluate the predictive validity of the value representations under different value systems. Using the Bradley-Terry model bradley1952rank, we train a linear classifier with pairwise cross-entropy loss to predict the relative safety of LLM pairs:
@@
[…]
@@ 5.3 LLM Value Alignment for Evaluating Representation Power
Values are cognitive representations of motivational goals schwartz2012overview. A well-constructed value system should be able to fully represent the broad motivations behind LLM outputs, therefore enabling effective transituational alignment of LLMs with human values.
Our evaluation builds on the BaseAlign algorithm yao2023value_fulcra, originally designed to align LLM outputs with Schwartz’s basic human values. This approach employs a target human value vector, either heuristically customized or derived from human survey data. We extend BaseAlign to arbitrary value systems by distilling the target vector from the human preference:
@@
[…]
@@
Next, we adopt PPO schulman2017proximal to align the LLM with the target value under $V$ yao2023value_fulcra:
[…]
where $\theta$ is the LLM policy parameters, $\mathcal{D}_{p}$ is the dataset of prompts dai2023safe, and $\pi_{\theta}(p)$ is the parameterized LLM response distribution. The representation power of $V$ is evaluated by the helpfulness and harmlessness of the aligned LLMs bai2022training; yao2023value_fulcra.
@@ 6 Results
This section presents the results of our experiments, including the proposed LLM value system, the analysis of the system, and the benchmarking of our system against the well-established Schwartz’s values schwartz2012overview.
@@ 6.1 Proposed Value System
Fig. 3 visualizes the value system constructed by GPLA. Table 1 gathers its factor loadings, Cronbach’s Alpha, and confidence intervals. The factor loadings indicate the strength of the relationship between each factor and its atomic values, with higher loadings suggesting more contribution to the factor. Cronbach’s Alpha measures the internal consistency of each factor, with higher values indicating greater reliability. Some atomic values are removed to ensure clear loading patterns and desirable factor reliability aavik2002structure. After that, all our factors reach the standard psychometric threshold of 0.7, indicating strong internal consistency taber2018use.
@@
[…]
@@ 6.2 Analyzing Value System
Fig. 3 (Top) presents the correlations between the factors in our value system. Similar to Schwartz’s theory of basic human values schwartz2012overview, LLM values also exhibit both compatible and opposing relationships. Notably, social responsibility, rule-following, and rationality show positive correlations with one another, while all of them are negatively correlated with risk-taking.
@@
Circumplex analysis is a statistical method that examines whether the underlying structure of variables aligns with a circumplex pattern, and, if so, the positions of variables on a circle. The stronger the correlation between variables, the shorter their distance on the circumference. We conduct circumplex analysis based on the correlations between factors. Fig. 3 (Bottom) illustrates the analysis results based on Browne’s circular stochastic process model browne1992circumplex; grassi2010circe. The compatible values are closer on the circle (e.g., Social Responsibility and Rule-Following) while opposing values are positioned diagonally (e.g., Risk-Taking and Rule-Following). The results verify the presence of a circumplex structure in the value system.
@@
To evaluate the consistency and robustness of our multivariate system structure (i.e., the 5-dimensional relationship), we measure LLM values using two distinct prompt datasets: the psychometric prompts from GPV ye2025gpv and the red-teaming prompts from SALAD-Bench li2024salad, which feature distinct prompt distributions. Their measurements yield an average intra-LLM correlation of 0.87; here we use intra-LLM correlations because the relative value hierarchy within an LLM is more important than their absolute measurements. This result indicates a high level of consistency in the value structure across prompt distributions. We also find a high correlation (0.73) between intra-LLM value consistency and LLM safety scores li2024salad. It suggests that LLMs with higher value consistency tend to be safer. Complete results are available in § B.3.
@@ 6.3 Comparing Value Systems
We benchmark the proposed value system against Schwartz’s value system schwartz2012overview, the most established framework for human values and commonly used in LLM value studies.
@@
We follow the standard validation procedures of CFA schwartz2004cfa to evaluate the structure validity of different value systems. For our value system, we construct it using half of the measurement data and bootstrap the data to ensure its sufficiency (#data points 5$\geq$ 5$\times$ 5 × #variables). The other half of the data is held out for CFA. For Schwartz’s value system, we map the observed value variables (the atomic values in our system) to its four high-level values or ten low-level values, according to the semantic relevance schwartz2004cfa measured by an embedding model openai2024textembedding3large; all data is used for CFA. Table 2 displays the CFA results. Our value system demonstrates a better fit for the data, capturing the underlying values behind LLM generations.
[…]
@@
[…]
Table 2: CFA results of different value systems. H and L denote high and low-level Schwartz values, respectively.
@@
We follow the experimental setup in (ye2025gpv, Section 5.2) to predict LLM safety based on their value orientations. § 6.3 presents the prediction accuracy under different value systems. The higher accuracy of our value system indicates its superior predictive validity. In addition, according to the parameters of well-trained linear classifiers, we find that Social Responsibility, Rule-Following, and Rationality enhance safety, whereas Risk-taking and Self-Competence undermine it; see § B.5 for details.
{wraptable}
[…]
@@
We follow the experimental setup in (yao2023value_fulcra, Section 6.2) to perform LLM value alignment. Different value systems are respectively used to represent LLM outputs and desired human values. We employ GPV ye2025gpv as an open-vocabulary value evaluator for all value systems, but also include the original results of yao2023value_fulcra using its Schwartz-specific evaluator for comparison. Table 3 shows the alignment performance of different value systems. Alignment under our value system converges to the lowest harmlessness and the highest helpfulness, establishing its superior representation power. Full experimental details are available in § A.4.
[…]
@@
[…]
Table 3: Alignment performance of different value systems. Schwartz* denotes the original results drawn from yao2023value_fulcra using a Schwartz-specific value evaluator. Both Schwartz baselines are based on the 10-dimensional Schwartz values yao2023value_fulcra.
@@ 7 Conclusion
This study presents GPLA, a novel psychologically grounded paradigm that enables automated, scalable, adaptable, and non-reactive value system construction. With GPLA, we introduce the first theoretically and empirically validated five-factor LLM value system, encompassing Social Responsibility, Risk-Taking, Rule-Following, Self-Competence, and Rationality. We accompany the value system with three benchmarking tasks, rooted in both psychological theories and AI priorities. Experimental results confirm the superior validity and utility of the proposed value system compared to the canonical Schwartz’s values.
GPLA promises to overcome the key limitations of traditional psycho-lexical paradigms by offering adaptability to evolving contexts and agent-specific needs. The proposed value system enables more effective evaluation, understanding, and alignment of LLM values, thereby contributing to their safe and reliable deployment.
@@
This study is limited to identifying and measuring values in English. Values are both culturally and linguistically sensitive schwartz2013culture, and LLM value orientations can vary across languages cahyawijaya2024high. Future work will extend GPLA to multilingual contexts.
In addition, we follow BaseAlign yao2023value_fulcra when benchmarking value systems on value alignment. However, the algorithm, like many existing alignment methods, assumes a single, consistent alignment target. Future work should explore distributional and pluralistic alignment sorensenposition; zhong2024panacea to more holistically evaluate a value system.
@@ A.1 Data Statistics for Collecting Value Lexicons
We collect value-laden LLM generations from four data sources: ValueBench ren2024valuebench, GPV ye2025gpv, ValueLex biedma2024beyond, and BeaverTails ji2024beavertails. They provide data of different forms: raw LLM responses, parsed perceptions (a sentence that is highly reflective of values ye2025gpv), and annotated values. The summary of the data statistics is shown in Table 4.
ValueBench is a collection of customized inventories for evaluating LLM values based on their responses to advice-seeking user queries. By administering the inventories to a set of LLMs, the authors collect 11,928 responses [^1] https://github.com/Value4AI/ValueBench/blob/main/assets/QA-dataset-answers-rating.xlsx, each considered as one perception. The responses are annotated with 37,526 values by Kaleido sorensen2024value, of which 330 are unique.
GPV ye2025gpv is a psychologically grounded framework for measuring LLM values given their free-form outputs. Perceptions are considered atomic measurement units in GPV, and the authors collect 24,179 perceptions [^2] https://github.com/Value4AI/gpv/blob/master/assets/question-answer-perception.csv from a set of LLM subjects. The perceptions are annotated with 62,762 values, of which 361 are unique.
In ValueLex biedma2024beyond, the authors collect 745 unique values from a set of fine-tuned LLMs via direct prompting (see Appendix C for more details).
BeaverTails ji2024beavertails is an AI safety-focused collection. We use a subset of the BeaverTails dataset [^3] https://huggingface.co/datasets/PKU-Alignment/BeaverTails/tree/main/round0/30k, which contains 3012 LLM responses, which are then parsed into 10,008 perceptions. The perceptions are annotated with 21,968 values, of which 395 are unique.
@@
[…]
@@ A.2 LLM Subjects
Our experiments involve 33 LLMs coupled with 21 profiling prompts rozen2024llms. The LLMs and profiling prompts are listed in Table 5 and Table 6, respectively.
[…]
@@
[…]
@@
[…]
Table 6: Anchoring prompts for LLM profiling rozen2024llms.
@@ A.3 Value Measurement for Structuring Value System
We measure the value orientations of the LLMs subjects following GPV ye2025gpv. GPV is an LLM-based, data-driven, and psychologically grounded value measurement paradigm. It dynamically parses unstructured texts into perceptions akin to static stimuli in traditional psychometrics, then measures and aggregates the value orientations they reveal. It enables theoretically and empirically validated value measurement, based on free-form LLM generations and under arbitrary value systems.
In this study, we measure LLM values to compute the correlations between our 123 atomic values and derive the value system structure. Similar to ye2025gpv, we generate one advice-seeking and value-eliciting prompt for each of the atomic values, using LABEL:prompt:item_generation. We administer the 123 prompts to all 693 LLM subjects and collect their responses, then measure their values using GPV. Each value dimension is measured using all responses. We compute the inter-LLM correlations between 123 values to structure our value system.
Prompt 1: Generating value-eliciting questions.
@@ A.4 LLM Value Alignment
All experiments were conducted on two NVIDIA L20 GPUs, each with 48GB of memory. We generally follow the experimental setup in yao2023value_fulcra, with the exceptions noted below. As shown in Table 3, our modifications improve the harmlessness of the aligned model with only marginal reduction in helpfulness.
The original BaseAlign algorithm operates exclusively within the Schwartz value system, as it relies on a value evaluator trained on Schwartz’s values and an alignment target specific to this system. We extend BaseAlign to align LLMs under any arbitrary value system. First, we employ GPV ye2025gpv as an open-vocabulary value evaluator. Second, we propose a method for distilling the alignment target from human preference data (§ 5.3). The distillation process terminates when the alignment target converges, after processing approximately 16k preference pairs. The results of this distillation are shown in Table 7 for Schwartz’s values and Table 8 for our system. The distilled alignment target, based on Schwartz’s values, closely matches the heuristically defined target in BaseAlign yao2023value_fulcra, demonstrating the effectiveness of our approach.
@@
[…]
@@
[…]
Table 7: Alignment targets for Schwartz’s values, on a scale from -1 to 1. Original: heuristically defined target in BaseAlign yao2023value_fulcra. Distilled: distilled target from human preference data.
[…]